# Optimal Baseline Corrections for Off-Policy Contextual Bandits

### v1

@@ Abstract.
submissionid: 92journalyear: 2024copyright: noneconference: Proceedings of the 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: Proceedings of the 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italy
@@ 1. Introduction & Motivation
Recommender systems have undergone a paradigm shift in the last few decades, moving their focus from rating prediction in the days of the Netflix Prize (Bennett et al., 2007), to item prediction from implicit feedback (Rendle, 2022) and ranking applications gaining practical importance (Steck, 2013). Recently, work that applies ideas from the algorithmic decision-making literature to recommendation problems has become more prominent (Vasile et al., 2020; Saito and Joachims, 2021; Gupta et al., 2024; Jeunen et al., 2022). While this line of research is not inherently new (Shani et al., 2002; Li et al., 2010), methods based on contextual bandits (or reinforcement learning by extension) have now become widespread in the recommendation field (McInerney et al., 2018; Mehrotra et al., 2020; Bendada et al., 2020; Jeunen and Goethals, 2021b; Yi et al., 2023; Su et al., 2024; Briand et al., 2024). The off-policy setting is particularly attractive for practitioners (van den Akker et al., 2024), as it allows models to be trained and evaluated in an offline manner (Chen et al., 2019; Dong et al., 2020; Ma et al., 2020; Jeunen et al., 2020; Jeunen and Goethals, 2021a; Chen et al., 2021; Liu et al., 2022; Chen et al., 2022; Jeunen and Goethals, 2023; Gupta et al., 2023b, a). Indeed, methods exist to obtain unbiased offline estimators of online reward metrics, which can then be optimized directly (Jeunen, 2021).
@@
In line with recent work on off-policy evaluation/learning for recommendation (Jeunen and Goethals, 2021b; Jeunen et al., 2023; Jeunen and Goethals, 2021a; Rohde et al., 2018; Saito et al., 2021), we adopt an off-policy simulation environment to emulate real-world recommendation scenarios, such as stochastic rewards, large action spaces, and controlled randomisation. This choice also encourages future reproducibility (Saito et al., 2020). Our experimental results indicate that baseline correction for gradient variance reduction enables substantially faster convergence and lower gradient variance during learning.
In addition, we derive a closed-form solution to the optimal baseline correction for off-policy evaluation, i.e., the one that minimizes the variance of the estimator itself. Importantly, since our framework only considers unbiased estimators, the variance-optimality implies overall optimality. Our experimental results show that this leads to lower error in policy value estimation, compared to widely used doubly-robust and SNIPS estimators (Dudík et al., 2014; Swaminathan and Joachims, 2015b).
@@ 2.1. On-policy contextual bandits
This provides an unbiased estimate of the gradient of $V(\pi_{\theta})$. However, it may be subject to high variance due to the inherent variance of $R$. Several techniques have been proposed in the literature that aim to alleviate this, mostly leveraging additive control variates.
Control variates are random variables with a known expectation (Owen, 2013, §8.9). If the control variate is correlated with the original estimand – in our case $V(\pi_{\theta})$ – they can be used to reduce the estimator’s variance. A natural way to apply control variates to a sample average estimate for Eq. 1 is to estimate a model of the reward $\widehat{r}(a,x)\approx\mathbb{E}[R|X=x;A=a]$ and subtract it from the observed rewards (Freedman, 2008). This is at the heart of key RL techniques (i.a., generalised advantage estimation (Schulman et al., 2016)), and it underpins widely used methods to increase sensitivity in online controlled experiments (Deng et al., 2013; Poyarkov et al., 2016; Budylin et al., 2018; Baweja et al., 2024). As such, it applies to both evaluation and learning tasks. We note that if the model $\widehat{r}(a,x)$ is biased, this bias propagates to the resulting estimator for $V(\pi_{\theta})$.
@@
This derivation follows the more general derivation from Greensmith et al. (2004) for partially observable Markov decision processes (POMDPs), we have not encountered its use in the existing bandit literature applied to recommendation problems. In section 3.2, we show that a similar line of reasoning can be applied to derive a variance-optimal gradient for the off-policy contextual bandit setup.
@@ 2.2. Off-policy estimation for general bandits
Importance sampling – the bread and butter of unbiased off-policy estimation – often leads to increased variance compared to on-policy estimators. Several variance reduction techniques have been proposed specifically to combat the excessive variance of $\widehat{V}_{\rm IPS}$ (Ionides, 2008; Dudík et al., 2014; Swaminathan and Joachims, 2015b). Within the scope of this work, we only consider techniques that reduce variance without introducing bias.
@@
Furthermore, as this random variable (Eq. 14) is likely to be correlated with the IPS estimates, we can expect that its use as a control variate will lead to reduced variance (see, e.g., (Kong, 1992)). This gives rise to the asymptotically unbiased and parameter-free self-normalised IPS (SNIPS) estimator, with $S\coloneqq\frac{1}{D}\sum_{(x,a,r)\in\mathcal{D}}\frac{\pi(a|x)}{\pi_{0}(a|x)}$ as its normalization term:
@@
(18) $\pi^{\star}=\operatorname*{arg\,max}_{\pi\in\Pi}\widehat{V}_{\rm IPS}(\pi,\mathcal{D}),\text{ s.t. }\frac{1}{\left|\mathcal{D}\right|}\sum_{(x,a,r)\in\mathcal{D}}\frac{\pi(a|x)}{\pi_{0}(a|x)}=S^{\star}.$
@@
Doubly robust estimation. Another way to reduce the variance of $\widehat{V}_{\rm IPS}$ is to leverage a model of the reward $\widehat{r}(a,x)\approx\mathbb{E}[R|X=x;A=a]$. Including it as an additive control variate in Eq. 12 gives rise to the doubly robust (DR) estimator, deriving its name from its unbiasedness if either the logging propensities $\pi_{0}$ or the reward model $\widehat{r}$ is unbiased (Dudík et al., 2014):
@@ 3.1. A unified off-policy estimator
From an optimization perspective, we are mainly interested the gradient of the $\widehat{V}_{{\rm\beta\text{-}IPS}}$ objective:
@@
Importantly, whilst these three approaches are motivated through entirely different lenses: minimizing gradient variance, applying a multiplicative control variate to reduce estimation variance, and applying an additive control variate to improve robustness. They result in equivalent gradients, and thus, in equivalent optima. Specifically, for optimization, the estimators are equivalent when $\beta\equiv\lambda\equiv\widehat{r}$.
This equivalence implies that the choice between these three approaches is not important. Since the simple baseline correction estimator $\widehat{V}_{\beta\text{-}{\rm IPS}}$ (Eq. 21) has an equivalence with all SNIPS estimators and all doubly-robust estimators with a constant reward. Thereby, we propose $\widehat{V}_{\beta\text{-}{\rm IPS}}$ should be seen as an estimator that unifies all three approaches. Accordingly, we argue that the real task is to find the optimal $\beta$ value for $\widehat{V}_{\beta\text{-}{\rm IPS}}$, since this results in an estimator that is at least as optimal as any estimator in the underlying families of estimators, and possibly superior to them.
@@ 3.2. Minimizing gradient variance
where we can ignore the the second term of the variance in Eq. 28, since it is independent of $\beta$ (Mohamed et al., 2020, Eq. 12). The optimal baseline can be obtained by solving for:
@@
Note that this expectation is over actions sampled by the logging policy. As a result, we can obtain Monte Carlo estimates of the corresponding expectations. The derivation has high similarity with the on-policy case (cf. Section 2.1(Greensmith et al., 2004). Nevertheless, we are unaware of any work on off-policy learning that leverages it. Joachims et al. (2018) refer to the on-policy variant with: “we cannot sample new roll-outs from the current policy under consideration, which means we cannot use the standard variance-optimal estimator used in REINFORCE.” Since the expectation is over actions sampled by the logging policy and not the target policy, we have shown that we do not need new roll-outs. Thereby, our estimation strategy is a novel off-policy approach that estimates the variance-optimal baseline.
@@ 3.3. Minimizing estimation variance
(39) ${\rm MSE}(\hat{V}(\pi))\coloneq\mathbb{E}_{\mathcal{D}}\left[(\hat{V}(\pi,,\mathcal{D})-V(\pi))^{2}\right].$
@@
(41) $\text{Bias}(\hat{V}(\pi),\mathcal{D})=|\mathbb{E}_{\mathcal{D}}\left[\hat{V}(\pi,\mathcal{D})-V(\pi,\mathcal{D})\right]|,$
@@ 1. RQ1
    Does the proposed estimator-variance-minimizing baseline correction (Eq. 38) improve OPL in a full-batch setting?
@@ 4. RQ4
Figure 3. Empirical variance of the gradient of different off-policy learning estimators in a mini-batch optimization setup with varying learning rates (from left to right). We compute the gradient variance for each mini-batch during the training and then report the average value across all mini-batches in a training epoch. Results are averaged across 32 independent runs with 95% confidence interval.
@@ 5.1. Off-policy learning performance (RQ1–3)
Full-batch. The results for the full-batch training in terms of the policy value on the test set are reported in Figure 1, over the number of training epochs. To minimize the impact of external factors, we use a linear model without bias, followed by a softmax to generate a distribution over all actions, given a context vector $x$ (this is a common setup, see, e.g., (Jeunen et al., 2020; Jeunen and Goethals, 2020; Sakhi et al., 2020)). We note that the goal of this work is not to get the maximum possible policy value on the test set but rather to evaluate the effect of baseline corrections on gradient and estimation variance. The simple model setup allows us to easily track the empirical gradient variance, given that we have only one parameter vector.
@@
Analogous to the full-batch setup, the IPS estimator results in a lower test policy value, most likely because of the high gradient variance which prevents convergence at high performance. In contrast, due to their baseline corrections, BanditNet (Eq. 24) and $\beta$-IPS have a lower gradient variance. Accordingly, they also converge at better performance (Bottou et al., 2018), i.e., resulting in superior test policy values.
Amongst these baseline-corrected gradient-based methods (BanditNet and $\beta$-IPS), our proposed $\beta$-IPS estimator outperforms BanditNet as it provides substantially higher policy. The differences are observed over different choices of learning rates. Thus we answer RQ2 accordingly: in the mini-batch setting, our proposed gradient-minimizing baseline method results in considerably higher policy value compared to both IPS and BanditNet.
Next, we directly consider the empirical gradient variance of different estimators; Figure 3 reports the average mini-batch gradient variance per epoch. As expected, the IPS estimator has the highest gradient variance by a large margin. For BanditNet, we observe a lower gradient variance, which is the desired result of the additive baseline it deploys. Finally, we observe that our proposed method $\beta$-IPS has the lowest gradient variance. This result corroborates the theoretical claim (Theorem 1), which states that the $\beta$-IPS has the lowest gradient variance amongst all global additive control variates (including IPS and BanditNet). Our answer to RQ3 is thus clear: our proposed $\beta$-IPS results in considerably lower gradient variance compared to BanditNet and IPS.
Figure 4. Mean Squared Error (MSE) of different off-policy estimators with varying action space (from left to right), and varying inverse temperature parameter of the softmax logging policy (from top to bottom). X-axis corresponds to the size of the logged data simulated (ranging from $10^{2}$ to $10^{6}$), and the y-axis corresponds to the MSE (evaluated over 100 independent samples of the synthetic data) along with 95% confidence interval. Each row corresponds to a different setting of inverse temp. of the softmax logging policy. We only consider unbiased (asymptotically or otherwise) estimators.
@@ 5.2. Off-policy evaluation performance (RQ4)
To evaluate the performance of the proposed $\beta$-IPS method, which minimizes the estimated policy value (Eq. 38), in an OPE task, results are presented in Figure 4. The target policy (to be evaluated) is a logistic regression model trained via the IPS objective on logged data and evaluated on a separate full-information test set. We evaluate the MSE of the estimated policy value against the true policy value (Eq. 39). To evaluate the MSE of different estimators realistically, we report results with varying degrees of the optimality of the behavior policy (decided by the inverse temperature parameter of the softmax) and with a varying cardinality of the action space. A positive (and higher) inverse softmax temp. results in a increasingly optimal behavior policy (selects action with highest probability), and a negative (and lower) inverse softmax temp. parameter results in an increasingly sub-optimal behavior policy (selects actions with lowest probability). Our proposed $\beta$-IPS method has the lowest MSE in all simulated settings. Interestingly, the proposed $\beta$-IPS has a lower MSE than the DR method, which has a regression model-based control variate, arguably more powerful than the constant control variate from the proposed $\beta$-IPS method. Similar observations have been made in previous work, e.g., Jeunen and Goethals (2020) reported that the DR estimator’s performance heavily depends on the randomization of the logging policy.
@@
Depending on the setting, we see that $\beta$-IPS either has performance comparable to the SNIPS estimator, i.e., when inverse temperature $\in\{-1,1\}$; or noticeably higher performance than SNIPS, i.e., when inverse temperature $\in\{-5,5\}$. We provide additional experimental results visualising the variance of the estimators (analogous to Figure 4) in the appendix (Section A).

### v2

@@ Abstract.
Contextual Bandits; Recommender Systems; Off-policy Learning
submissionid: 92journalyear: 2024copyright: rightsretainedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688105isbn: 979-8-4007-0505-2/24/10ccs: Information systems Recommender systems
@@ 1. Introduction & Motivation
Recommender systems have undergone a paradigm shift in the last few decades, moving their focus from rating prediction in the days of the Netflix Prize (Bennett et al., 2007), to item prediction from implicit feedback (Rendle, 2022) and ranking applications gaining practical importance (Steck, 2013; Jeunen et al., 2024). Recently, work that applies ideas from the algorithmic decision-making literature to recommendation problems has become more prominent (Vasile et al., 2020; Saito and Joachims, 2021; Gupta et al., 2024a; Jeunen et al., 2022). While this line of research is not inherently new (Shani et al., 2002; Li et al., 2010), methods based on contextual bandits (or reinforcement learning by extension) have now become widespread in the recommendation field (McInerney et al., 2018; Mehrotra et al., 2020; Bendada et al., 2020; Jeunen and Goethals, 2021b; Yi et al., 2023; Su et al., 2024; Briand et al., 2024). The off-policy setting is particularly attractive for practitioners (van den Akker et al., 2024), as it allows models to be trained and evaluated in an offline manner (Chen et al., 2019; Dong et al., 2020; Ma et al., 2020; Jeunen et al., 2020; Jeunen and Goethals, 2021a; Chen et al., 2021; Liu et al., 2022; Chen et al., 2022; Jeunen and Goethals, 2023; Gupta et al., 2023c, a, 2024b, b). Indeed, methods exist to obtain unbiased offline estimators of online reward metrics, which can then be optimized directly (Jeunen, 2021).
@@
In line with recent work on off-policy evaluation/learning for recommendation (Jeunen and Goethals, 2021b; Jeunen et al., 2023; Jeunen and Goethals, 2021a; Rohde et al., 2018; Saito et al., 2021), we adopt an off-policy simulation environment to emulate real-world recommendation scenarios, such as stochastic rewards, large action spaces, and controlled randomisation. This choice also encourages future reproducibility (Saito et al., 2020). Our experimental results indicate that our proposed baseline correction for gradient variance reduction enables substantially faster convergence and lower gradient variance during learning.
In addition, we derive a closed-form solution to the optimal baseline correction for off-policy evaluation, i.e., the one that minimizes the variance of the estimator itself. Importantly, since our framework only considers unbiased estimators, the variance-optimality implies overall optimality. Our experimental results show that this leads to lower errors in policy value estimation than widely used doubly-robust and SNIPS estimators (Dudík et al., 2014; Swaminathan and Joachims, 2015b).
All source code to reproduce our experimental results is available at: https://github.com/shashankg7/recsys2024_optimal_baseline.
@@ 2.1. On-policy contextual bandits
This provides an unbiased estimate of the gradient of $V(\pi_{\theta})$. However, it may be subject to high variance due to the inherent variance of $R$. Several techniques have been proposed in the literature that aim to alleviate this, mostly using additive control variates.
Control variates are random variables with a known expectation (Owen, 2013, §8.9). If the control variate is correlated with the original estimand — in our case $V(\pi_{\theta})$ — they can be used to reduce the estimator’s variance. A natural way to apply control variates to a sample average estimate for Eq. 1 is to estimate a model of the reward $\widehat{r}(a,x)\approx\mathbb{E}[R|X=x;A=a]$ and subtract it from the observed rewards (Freedman, 2008). This is at the heart of key RL techniques (i.a., generalised advantage estimation (Schulman et al., 2016)), and it underpins widely used methods to increase sensitivity in online controlled experiments (Deng et al., 2013; Poyarkov et al., 2016; Budylin et al., 2018; Baweja et al., 2024). As such, it applies to both evaluation and learning tasks. We note that if the model $\widehat{r}(a,x)$ is biased, this bias propagates to the resulting estimator for $V(\pi_{\theta})$.
@@
This derivation follows the more general derivation from Greensmith et al. (2004) for partially observable Markov decision processes (POMDPs). We have not encountered its use in the existing bandit literature applied to recommendation problems. In Section 3.2, we show that a similar line of reasoning can be applied to derive a variance-optimal gradient for the off-policy contextual bandit setup.
@@ 2.2. Off-policy estimation for general bandits
Importance sampling — the bread and butter of unbiased off-policy estimation — often leads to increased variance compared to on-policy estimators. Several variance reduction techniques have been proposed specifically to combat the excessive variance of $\widehat{V}_{\rm IPS}$ (Ionides, 2008; Dudík et al., 2014; Swaminathan and Joachims, 2015b). Within the scope of this work, we only consider techniques that reduce variance without introducing bias.
@@
Furthermore, as this random variable (Eq. 14) is likely to be correlated with the IPS estimates, we can expect that its use as a control variate will lead to reduced variance (see (e.g., Kong, 1992)). This gives rise to the asymptotically unbiased and parameter-free self-normalised IPS (SNIPS) estimator, with $S\coloneqq\frac{1}{D}\sum_{(x,a,r)\in\mathcal{D}}\frac{\pi(a|x)}{\pi_{0}(a|x)}$ as its normalization term:
@@
(18) $\pi^{\star}=\operatorname*{arg\,max}_{\pi\in\Pi}\widehat{V}_{\rm IPS}(\pi,\mathcal{D}),\text{ s.th. }\frac{1}{\left|\mathcal{D}\right|}\sum_{(x,a,r)\in\mathcal{D}}\frac{\pi(a|x)}{\pi_{0}(a|x)}=S^{\star}.$
@@
Doubly robust estimation. Another way to reduce the variance of $\widehat{V}_{\rm IPS}$ is to use a model of the reward $\widehat{r}(a,x)\approx\mathbb{E}[R|X=x;A=a]$. Including it as an additive control variate in Eq. 12 gives rise to the doubly robust (DR) estimator, deriving its name from its unbiasedness if either the logging propensities $\pi_{0}$ or the reward model $\widehat{r}$ is unbiased (Dudík et al., 2014):
@@ 3.1. A unified off-policy estimator
From an optimization perspective, we are mainly interested in the gradient of the $\widehat{V}_{{\rm\beta\text{-}IPS}}$ objective:
@@
Importantly, these three approaches are motivated through entirely different lenses: minimizing gradient variance, applying a multiplicative control variate to reduce estimation variance, and applying an additive control variate to improve robustness. But they result in equivalent gradients, and thus, in equivalent optima. Specifically, for optimization, the estimators are equivalent when $\beta\equiv\lambda\equiv\widehat{r}$.
This equivalence implies that the choice between these three approaches is not important. Since the simple baseline correction estimator $\widehat{V}_{\beta\text{-}{\rm IPS}}$ (Eq. 21) has an equivalence with all SNIPS estimators and all doubly-robust estimators with a constant reward, we propose that $\widehat{V}_{\beta\text{-}{\rm IPS}}$ should be seen as an estimator that unifies all three approaches. Accordingly, we argue that the real task is to find the optimal $\beta$ value for $\widehat{V}_{\beta\text{-}{\rm IPS}}$, since this results in an estimator that is at least as optimal as any estimator in the underlying families of estimators, and possibly superior to them.
@@ 3.2. Minimizing gradient variance
where we can ignore the second term of the variance in Eq. 28, since it is independent of $\beta$ (Mohamed et al., 2020, Eq. 12). The optimal baseline can be obtained by solving for:
@@
Note that this expectation is over actions sampled by the logging policy. As a result, we can obtain Monte Carlo estimates of the corresponding expectations. The derivation has high similarity with the on-policy case (cf. Section 2.1(Greensmith et al., 2004). Nevertheless, we are unaware of any work on off-policy learning that uses it. Joachims et al. (2018) refer to the on-policy variant with: “we cannot sample new roll-outs from the current policy under consideration, which means we cannot use the standard variance-optimal estimator used in REINFORCE.” Since the expectation is over actions sampled by the logging policy and not the target policy, we have shown that we do not need new roll-outs. Thereby, our estimation strategy is a novel off-policy approach that estimates the variance-optimal baseline.
@@ 3.3. Minimizing estimation variance
(39) ${\rm MSE}(\hat{V}(\pi))\coloneq\mathbb{E}_{\mathcal{D}}\left[(\hat{V}(\pi,\mathcal{D})-V(\pi))^{2}\right].$
@@
(41) $\text{Bias}(\hat{V}(\pi),\mathcal{D})=\left|\mathbb{E}_{\mathcal{D}}\left[\hat{V}(\pi,\mathcal{D})-V(\pi)\right]\right|,$
@@ 1. RQ1
    Does the proposed estimator-variance-minimizing baseline correction (Eq. 38) improve off-policy learning (OPL) in a full-batch setting?
@@ 4. RQ4
Figure 3. Empirical variance of the gradient of different off-policy learning estimators in a mini-batch optimization setup with varying learning rates (in title). We compute gradient variance for each mini-batch during training and then report the average value across all mini-batches in a training epoch. Results are averaged across 32 independent runs with 95% confidence interval.
@@ 5.1. Off-policy learning performance (RQ1–3)
Full-batch. The results for the full-batch training in terms of the policy value on the test set are reported in Figure 1, over the number of training epochs. To minimize the impact of external factors, we use a linear model without bias, followed by a softmax to generate a distribution over all actions, given a context vector $x$ (this is a common setup, see (e.g., Jeunen et al., 2020; Jeunen and Goethals, 2020; Sakhi et al., 2020)). We note that the goal of this work is not to get the maximum possible policy value on the test set but rather to evaluate the effect of baseline corrections on gradient and estimation variance. The simple model setup allows us to easily track the empirical gradient variance, given that we have only one parameter vector.
@@
Analogous to the full-batch setup, the IPS estimator results in a lower test policy value, most likely because of the high gradient variance which prevents convergence to high performance. In contrast, due to their baseline corrections, BanditNet (Eq. 24) and $\beta$-IPS have a lower gradient variance. Accordingly, they also converge to better performance (Bottou et al., 2018), i.e., resulting in superior test policy values.
Amongst these baseline-corrected gradient-based methods (BanditNet and $\beta$-IPS), our proposed $\beta$-IPS estimator outperforms BanditNet as it provides a policy with substantially higher value. The differences are observed over different choices of learning rates. Thus we answer RQ2 accordingly: in the mini-batch setting, our proposed gradient-minimizing baseline method results in considerably higher policy value compared to both IPS and BanditNet.
Figure 4. Mean Squared Error (MSE) of different off-policy estimators with varying action space (from left to right), and varying inverse temperature parameter of the softmax logging policy (from top to bottom). X-axis corresponds to the size of the logged data simulated (ranging from $10^{2}$ to $10^{6}$), and the y-axis corresponds to the MSE (evaluated over 100 independent samples of the synthetic data) along with 95% confidence interval. Each row corresponds to a different setting of inverse temperature of the softmax logging policy. We only consider unbiased (asymptotically or otherwise) estimators.
Next, we directly consider the empirical gradient variance of different estimators; Figure 3 reports the average mini-batch gradient variance per epoch. As expected, the IPS estimator has the highest gradient variance by a large margin. For BanditNet, we observe a lower gradient variance, which is the desired result of the additive baseline it employs. Finally, we observe that our proposed method $\beta$-IPS has the lowest gradient variance. This result corroborates the theoretical claim (Theorem 1) that the $\beta$-IPS estimator has the lowest gradient variance amongst all global additive control variates (including IPS and BanditNet). Our answer to RQ3 is thus clear: our proposed $\beta$-IPS results in considerably lower gradient variance compared to BanditNet and IPS.
@@ 5.2. Off-policy evaluation performance (RQ4)
To evaluate the performance of the proposed $\beta$-IPS method, which minimizes the estimated policy value (Eq. 38), in an OPE task, results are presented in Figure 4. The target policy (to be evaluated) is a logistic regression model trained via the IPS objective on logged data and evaluated on a separate full-information test set. We evaluate the MSE of the estimated policy value against the true policy value (Eq. 39). To evaluate the MSE of different estimators realistically, we report results with varying degrees of the optimality of the behavior policy (decided by the inverse temperature parameter of the softmax) and with a varying cardinality of the action space. A positive (and higher) inverse softmax temperature results in a increasingly optimal behavior policy (selects action with highest reward probability), and a negative (and lower) inverse softmax temperature parameter results in an increasingly sub-optimal behavior policy (selects actions with lowest reward probability). Our proposed $\beta$-IPS method has the lowest MSE in all simulated settings. Interestingly, the proposed $\beta$-IPS has a lower MSE than the DR method, which has a regression model-based control variate, arguably more powerful than the constant control variate from the proposed $\beta$-IPS method. Similar observations have been made in previous work, e.g., Jeunen and Goethals (2020) reported that the DR estimator’s performance heavily depends on the logging policy.
@@
Depending on the setting, we see that $\beta$-IPS either has performance comparable to the SNIPS estimator, i.e., when inverse temperature $\in\{-1,1\}$; or noticeably higher performance than SNIPS, i.e., when inverse temperature $\in\{-5,5\}$.
@@
Figure 5. Empirical variance of different off-policy estimators with varying action space (from left to right), and varying sub-optimality of a temperature-based softmax behavior policy (from top to bottom). The x-axis corresponds to the size of the logged data simulated (ranging from $10^{2}$ to $10^{6}$), and the y-axis corresponds to the variance of different estimators (evaluated over 100 independent samples of the synthetic data) along with 95% confidence interval. Each row corresponds to a different optimality level of the logging policy, decided by the inverse temperature parameter. We only consider unbiased (asymptotically or otherwise) estimators.